# Blind and Topological Interference Managements for Bistatic Integrated Sensing and Communication

### v4

@@ V-A Proof of Theorem 4-a: Topology 1 with Non-cooperative Transmitters
The interference dimension from communication signals is the same as that for communication users, i.e., ${\mathsf{K}}-D$. Therefore, we can get a left null vector ${\mathbf{v}}_{0}$ in $\begin{bmatrix}{\mathbf{v}}^{[D+1]}{\mathbf{v}}^{[D+2]}\ldots{\mathbf{v}}^{[{\mathsf{K}}]}\end{bmatrix}$ to remove the communication interference within ${\mathsf{K}}-D+U+1$ time slots, thus
${\mathbf{v}}_{0}{\mathbf{y}}^{[{\mathsf{K}}+1]}$
$=H^{[{\mathsf{K}}+1,{\mathsf{K}}+1]}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}\negmedspace+\negmedspace\sum_{i=D+1}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0}{\mathbf{x}}^{[i]}\negmedspace+\negmedspace{\mathbf{v}}_{0}{\mathbf{z}}^{[{\mathsf{K}}+1]}$
$=H^{[{\mathsf{K}}+1,{\mathsf{K}}+1]}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}+\sum_{i=D+1}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[i]}+z^{\prime[{\mathsf{K}}+1]}.$
Let $x_{1}^{[0]}=\begin{bmatrix}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}&{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[D+1]}&\cdots&{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[K-U]}\end{bmatrix}^{H}$, in other $({\mathsf{K}}-D-U)$ transmission, we can use the same way to get $x_{2}^{[0]},x_{3}^{[0]},\cdots,x_{{\mathsf{K}}-D-U+1}^{[0]}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U+1}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},\cdots,x_{{\mathsf{K}}-D-U+1}^{[0]}]^{H}=I$. So one effective observation of the channel can be obtained, achieving the sDoF equal to $\frac{1}{{\mathsf{K}}-D+U+1}$.
V-B Proof of Theorem 4-b: Topology 2 with Non-cooperative Transmitters
Considering the bistatic ISAC systems with heterogeneous connectivity and non-cooperative transmitters, we describe the main idea of the proposed TIM scheme for the Topology 2 through the following example, while the general description can be found in the Appendix B and Appendix C.
Figure 7: Non-cooperative interference channel in Example 3.
Example 3 (Topology 2 with non-cooperative transmitters).
Consider the bistatic ISAC system with partially connected interference channel including $5$ single-antenna transmitters, $5$ single-antenna receivers and one sensor, as illustrated in Fig. 7. Each receiver or sensor is disconnected from the previous $U=1$ and next $D=1$ transmitters, and connected to other $3$ transmitters; the topology belongs to Topology 2 from the $(5,1,1)$-neighboring antidotes network described in Definition 2. We have
${\mathcal{R}}^{[1]}_{c}=\{1,3,4\},\ {\mathcal{R}}^{[2]}_{c}=\{2,4,5\},\ {\mathcal{R}}^{[3]}_{c}=\{1,3,5\},$
${\mathcal{R}}^{[4]}_{c}=\{1,2,4\},\ {\mathcal{R}}^{[5]}_{c}=\{2,3,5\},\ {\mathcal{R}}_{s}=\{3,4\}.$
For this system, the tradeoff points $(\text{sDoF},\text{cDoF})=(1,0)$, $(\text{sDoF},\text{cDoF})=(0,2)$ can be achieved by the sensing-only scheme and communication-only scheme, respectively. We then propose a TIM ISAC scheme with $(\text{sDoF},\text{cDoF})=(1/5,2)$.
Let $t_{0}=10$. During these $10$ time slots, the sensor will obtain $2$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}\in\mathbb{C}^{3\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}]^{H}=I$. In each period of $5$ time slots, we let each receiver recover $2$ communication messages, and let the sensor obtain one effective observation. In the following, we illustrate our proposed scheme for the first $5$ time slots, where each receiver $j\in\{1,\ldots,5\}$ should recover $2$ communication messages $W_{1}^{[j]},W_{2}^{[j]}\in\mathbb{C}$.
Denote the transmission signals of transmitter $j$ by ${\mathbf{x}}^{[j]}=[{\mathbf{x}}^{[j]}(1),\ldots,{\mathbf{x}}^{[j]}(5)]^{T}\in\mathbb{C}^{5\times 1}$. Then for the transmitters not connected to the sensor, we design
${\mathbf{x}}^{[1]}={\mathbf{v}}^{[1]}W_{1}^{[1]}+{\mathbf{v}}^{[2]}W_{2}^{[1]},$
${\mathbf{x}}^{[2]}={\mathbf{v}}^{[2]}W_{1}^{[2]}+{\mathbf{v}}^{[3]}W_{2}^{[2]},$
${\mathbf{x}}^{[5]}={\mathbf{v}}^{[5]}W_{1}^{[5]}+{\mathbf{v}}^{[1]}W_{2}^{[5]};$
for the transmitters connected to the sensor, we design
${\mathbf{x}}^{[4]}={\mathbf{v}}^{[4]}W_{1}^{[4]}+{\mathbf{v}}^{[5]}W_{2}^{[4]}+{\mathbf{x}}_{s}^{[4]}.$
Note that ${\mathbf{v}}^{[1]},\ldots,{\mathbf{v}}^{[5]}\in\mathbb{C}^{5\times 1}$ represent $5$ random i.i.d. vectors, which are linearly independent, and ${\mathbf{x}}_{s}^{[3]},{\mathbf{x}}_{s}^{[4]}\in\mathbb{C}^{5\times 1}$ are the transmitted sensing signals to be determined later.
In this partially connected case, by exploiting the network topology, we align the interference to each receiver. Let us consider receiver $1$ as an example. The received signal after removing the dedicated sensing signals at time slot $t\in\{1,\ldots,5\}$ is $\hat{y}^{[1]}(t)=\sum_{i\in\mathcal{R}_{1}=\{1,3,4\}}H^{[1i]}(t){\mathbf{x}}^{[i]}(t)+z^{[1]}(t).$ Hence, its received signals from the $5$ time slots can be formulated as:
$\hat{{\mathbf{y}}}^{[1]}=\begin{bmatrix}y^{[1]}(1)\\ \vdots\\ y^{[1]}(5)\\ \end{bmatrix}=H^{[11]}\left({\mathbf{v}}^{[1]}W_{1}^{[1]}+{\mathbf{v}}^{[2]}W_{2}^{[1]}\right)$
$\quad+H^{[13]}\left({\mathbf{v}}^{[3]}W_{1}^{[3]}+{\mathbf{v}}^{[4]}W_{2}^{[3]}\right)$
$\quad=\underbrace{{\mathbf{v}}^{[1]}H^{[11]}W_{1}^{[1]}+{\mathbf{v}}^{[2]}H^{[11]}W_{2}^{[1]}}_{\text{desired signals}}+{\mathbf{v}}^{[3]}H^{[13]}W_{1}^{[3]}$
$\quad+{\mathbf{v}}^{[4]}\left(H^{[13]}W_{2}^{[3]}+H^{[14]}W_{1}^{[4]}\right)+{\mathbf{v}}^{[5]}H^{[14]}W_{2}^{[4]}+{\mathbf{z}}^{[1]}.$
Thus in $\hat{{\mathbf{y}}}^{[1]}$ the dimension of the designed signals is $2$, and the dimension of the interference is aligned to $3$. By the linearly independence of ${\mathbf{v}}^{[1]},{\mathbf{v}}^{[2]},{\mathbf{v}}^{[3]},{\mathbf{v}}^{[4]},{\mathbf{v}}^{[5]}$, receiver $1$ can recover $W_{1}^{[1]},W_{2}^{[1]}$ from the first $5$ time slots. Similarly, each receiver can recover $2$ communication messages from the first $5$ time slots; thus the cDoF is $2$.
For the sensor, its received signal at time slot $t\in\{1,\ldots,5\}$ is $y^{[6]}(t)=\sum_{i\in\mathcal{R}_{s}=\{3,4\}}H^{[6i]}(t){\mathbf{x}}^{[i]}(t)+z^{[6]}(t).$ Hence, its received signals from the $5$ time slots can be formulated as:
$+H^{[64]}\left({\mathbf{v}}^{[4]}W_{1}^{[4]}+{\mathbf{v}}^{[5]}W_{2}^{[4]}+{\mathbf{x}}_{s}^{[4]}\right)+{\mathbf{z}}^{[6]}$
$+{\mathbf{v}}^{[5]}H^{[64]}W_{2}^{[4]}+H^{[53]}{\mathbf{x}}_{s}^{[3]}+H^{[54]}{\mathbf{x}}_{s}^{[4]}+{\mathbf{z}}^{[6]}.$
Note that $[{\mathbf{v}}^{[3]},{\mathbf{v}}^{4]},{\mathbf{v}}^{[5]}]$ with dimension $5\times 3$ has $2$ linearly independent left null vectors with high probability; let ${\mathbf{v}}_{0}\in\mathbb{C}^{1\times 5}$ be a non-zero left null vector. Thus we have
${\mathbf{v}}_{0}{\mathbf{y}}^{[6]}={\mathbf{v}}_{0}{\mathbf{v}}^{[3]}H^{[53]}W_{1}^{[3]}+{\mathbf{v}}_{0}{\mathbf{v}}^{[4]}\left(H^{[63]}W_{2}^{[3]}+H^{[64]}W_{1}^{[4]}\right)$
$+{\mathbf{v}}_{0}{\mathbf{v}}^{[5]}H^{[54]}W_{2}^{[4]}+{\mathbf{v}}_{0}H^{[53]}{\mathbf{x}}_{s}^{[3]}+{\mathbf{v}}_{0}H^{[54]}{\mathbf{x}}_{s}^{[4]}+{\mathbf{v}}_{0}{\mathbf{z}}^{[5]}$
$=\begin{bmatrix}H^{[53]}&H^{[54]}\end{bmatrix}\begin{bmatrix}{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]}\\ {\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[4]}\end{bmatrix}+{\mathbf{z}}^{\prime[5]},$
where $z^{\prime[5]}={\mathbf{v}}_{0}{\mathbf{z}}^{[5]}$. Hence, the sensor can eliminate the interference from communication messages, and obtain an effective observation from the first $5$ time slots.
We then fix the values of ${\mathbf{x}}_{s}^{[3]},{\mathbf{x}}_{s}^{[4]}$. Recall that ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}$ has been selected before, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]}]^{H}=I$. Now we let $x_{1}^{[0]}=[{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]},{\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[4]}]$; in other words, we solve ${\mathbf{x}}_{s}^{[3]}$ with $5$ parameters satisfying ${\mathbf{v}}_{0}{\mathbf{x}}_{s}^{[3]}=x_{1,1}$, where $x_{1,1}$ is the first element of ${\mathbf{x}}_{1}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[4]}$. Therefore, the sensing degree of freedom is $1/5$.
V-C Proof of Theorem 5-a: Topology 1 with Cooperative Transmitters
@@
We then consider the bistatic ISAC systems with heterogeneous connectivity and cooperative transmitters. For the Topology 1 based on the $({\mathsf{K}}+1,d)$-regular network, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{1}{d+1},\frac{2{\mathsf{K}}}{d+1}\right)$. Recall that the network topology could be expressed as
@@
Denote the transmitted signals of transmitter $j\in\{1,\ldots,{\mathsf{K}}+1\}$ by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}(d+1)]^{T}\in\mathbb{C}^{(d+1)\times 1}$. For the $d$ transmitters connected to the sensor, let
@@
For the transmitters in the set $\{d,d+1,\ldots,{\mathsf{K}}\}$, which are not connected to the sensor, let
@@
The $k\in\{d,\ldots,{\mathsf{K}}-d+1\}$-th receiver only connected to the communication transmitter, the received signal is:
@@
The dimension of the desired signals is 2, and the dimension of the interference signals is $d-1$, we can achieve symmetric $\text{cDoF}=\frac{2}{d+1}$ for each receiver. Note that all interference at the receivers originates from communication signals, while sensing signals can be eliminated. The receiver that only receives communication signals experiences the strongest interference. For receivers that acquire both sensing and communication signals, the interference dimension does not exceed $d-1$. Given that the useful signal decoded by all receivers is 2, the communication degrees of freedom for all other receivers is $\frac{2}{d+1}$ as well. Thus, the cDoF is $\frac{2{\mathsf{K}}}{d+1}$.
For the sensor, the received signal ${\mathbf{y}}^{\left[s\right]}$ is:
@@
$\quad+H^{[{\mathsf{K}}+1,d-1]}\left({\mathbf{x}}_{s}^{[d-1]}+{\mathbf{v}}^{[d-1]}W_{1}^{[d-1]}\right)$
$=\sum^{d-1}_{j=1}{\mathbf{v}}^{[j]}(H^{[{\mathsf{K}}+1,j]}W_{1}^{[j]}+H^{[{\mathsf{K}}+1,j-1]}W_{2}^{[j-d]})+{\mathbf{z}}^{[{\mathsf{K}}+1]}.$
Note that $[{\mathbf{v}}^{\left[1\right]}{\mathbf{v}}^{\left[2\right]}\cdots{\mathbf{v}}^{\left[d-1\right]}]$ with dimension $(d+1)\times(d-1)$ has 2 linearly independent left null vectors with high probability; let ${\mathbf{v}}_{0}\ in{\mathbb{C}}^{1\times d+1}$ be a non-zero left null vector. Thus
$v_{0}{\mathbf{y}}^{\left[s\right]}\negmedspace=\negmedspace\sum_{j=1}^{d-1}\negmedspace\negmedspace H^{\left[K+1,j\right]}v_{0}{\mathbf{x}}_{s}^{\left[j\right]}\negmedspace+\negmedspace H^{\left[K+1,K+1\right]}v_{0}{\mathbf{x}}_{s}^{\left[K+1\right]}\negmedspace+\negmedspace{\mathbf{v}}_{0}{\mathbf{z}}^{[{\mathsf{K}}+1]}.$
Let $x_{1}^{\left[0\right]}=\left[v_{0}{\mathbf{x}}_{s}^{\left[1\right]},v_{0}{\mathbf{x}}_{s}^{\left[2\right]},\cdots,v_{0}{\mathbf{x}}_{s}^{\left[d-1\right]},v_{0}{\mathbf{x}}_{s}^{\left[{\mathsf{K}}+1\right]}\right]$, and this gives one valid observation. So one effective observation of the channel can be obtained, achieving a sDoF of $\frac{1}{d+1}$.

### v5

@@ V-A Proof of Theorem 4: $({\mathsf{K}}+1,U,D)$ neighboring antidotes network
The interference dimension from communication signals is the same as that for receivers, i.e., ${\mathsf{K}}-D$. Therefore, we can get $U$ linearly independent left null vectors ${\mathbf{v}}_{0,1},{\mathbf{v}}_{0,2},\dots,{\mathbf{v}}_{0,U}$ of $\begin{bmatrix}{\mathbf{v}}^{[D+1]},{\mathbf{v}}^{[D+2]},\ldots,{\mathbf{v}}^{[{\mathsf{K}}]}\end{bmatrix}$ to remove the communication interference within ${\mathsf{K}}-D+U+1$ time slots. Thus
${\mathbf{v}}_{0,u}{\mathbf{y}}^{[{\mathsf{K}}+1]}$ $=H^{[{\mathsf{K}}+1,{\mathsf{K}}+1]}{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}$
$+\sum_{i=D+1}^{{\mathsf{K}}-U}H^{[{\mathsf{K}}+1,i]}{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[i]}+{\mathbf{v}}_{0,u}{\mathbf{z}}^{[{\mathsf{K}}+1]}.$
We then fix the values of ${\mathbf{x}}_{s}^{[{\mathsf{K}}+1]},{\mathbf{x}}_{s}^{[D+1]},\dots,{\mathbf{x}}_{s}^{[{\mathsf{K}}-U]}$, the orthogonal sensing signals ${\mathbf{x}}_{u}^{[0]},u\in\{1,2,\dots,U+1\}$ have been generated before. Let ${\mathbf{x}}_{u}^{[0]}=\begin{bmatrix}{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]},{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[D+1]},\ldots,{\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[{\mathsf{K}}-U]}\end{bmatrix}^{T}$, we solve ${\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}$ with ${\mathsf{K}}-D+U+1$ elements satisfying ${\mathbf{v}}_{0,u}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}=x_{u,1}$, where $x_{u,1}$ is the first element of ${\mathbf{x}}_{u}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[D+1]},\dots,{\mathbf{x}}_{s}^{[{\mathsf{K}}-U]}$. $U+1$ effective observations of the channel can be acquired; all transmission can achieve an sDoF equal to $\frac{U+1}{{\mathsf{K}}-D+U+1}$.
V-B Proof of Theorem 5: $({\mathsf{K}}+1,d)$ regular network
@@
We then consider a bistatic ISAC system with heterogeneous connectivity, modeled as a $({\mathsf{K}}+1,d)$-regular network. For this network, we propose an ISAC scheme based on TIM achieving the tradeoff point $(\text{sDoF},\text{cDoF})=\left(\frac{2}{d+1},\frac{2{\mathsf{K}}}{d+1}\right)$. Recall that the network topology could be expressed as
@@
Let $t_{0}=(d+1)\left\lceil\frac{d}{2}\right\rceil$. In $(d+1)\left\lceil\frac{d}{2}\right\rceil$ time slots, the sensor will obtain $N=2\left\lceil\frac{d}{2}\right\rceil$ effective observations, through the sensing signals ${\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}\in{\mathbb{C}}^{(d+1)\times 1}$, satisfying $[{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}][{\mathbf{x}}_{1}^{[0]},{\mathbf{x}}_{2}^{[0]},...,{\mathbf{x}}_{N}^{[0]}]^{H}=\textbf{I}_{(d+1)}$. In each period of $d+1$ time slots, we let each receiver decode $2$ communication message symbols, and let the sensor obtain $2$ effective observations. Next we will illustrate our proposed scheme for the first $d+1$ time slots, where each receiver $k\in\{1,2,...,{\mathsf{K}}\}$ should recover $2$ communication message symbols $W_{1}^{[k]},W_{2}^{[k]}\in{\mathbb{C}}$.
Denote the transmitted signals of $j$-th transmitter, $j\in\{1,\ldots,{\mathsf{K}}+1\}$ by ${\mathbf{x}}^{[j]}=[x^{[j]}(1),\ldots,x^{[j]}(d+1)]^{T}\in\mathbb{C}^{(d+1)\times 1}$. For the $d$ transmitters connected to the sensor, let
@@
For the transmitters in $\{d,d+1,\ldots,{\mathsf{K}}\}$, which are not connected to the sensor, let
@@
For receiver $k\in\{d,\ldots,{\mathsf{K}}-d+1\}$ which is only connected to the communication transmitter, the received signal is:
@@
The dimension of the desired signals is 2, and the dimension of the interference signals is $d-1$, we can achieve symmetric $\text{cDoF}=\frac{2}{d+1}$ for each receiver. Note that all interference at the receivers originates from communication signals, while sensing signals can be eliminated. The receiver that only receives communication signals experiences the strongest interference. For the receivers that receive both sensing and communication signals (the receivers not in $\{d,\ldots,{\mathsf{K}}-d+1\}$), the interference dimension does not exceed $d-1$; thus in $d+1$ time slots, each of them can also decode $2$ communication message symbols, achieving the cDoF $\frac{2}{d+1}$ per receiver. Thus, the sum cDoF of the whole system is $\frac{2{\mathsf{K}}}{d+1}$.
For the sensor, the received signal is:
@@
$\quad+H^{[{\mathsf{K}}+1,d-1]}\left({\mathbf{x}}_{s}^{[d-1]}+{\mathbf{v}}^{[d-1]}W_{1}^{[d-1]}\right)+{\mathbf{z}}^{[{\mathsf{K}}+1]}$
$=\sum^{d-1}_{j=1}{\mathbf{v}}^{[j]}(H^{[{\mathsf{K}}+1,j]}W_{1}^{[j]}+H^{[{\mathsf{K}}+1,j-1]}W_{2}^{[j-d]})$
$\quad+\sum_{j=1}^{d}H^{[{\mathsf{K}}+1,j-1]}{\mathbf{x}}_{s}^{[j-1]}+{\mathbf{z}}^{[{\mathsf{K}}+1]}.$
Note that $[{\mathbf{v}}^{\left[1\right]},{\mathbf{v}}^{\left[2\right]},\ldots,{\mathbf{v}}^{\left[d-1\right]}]$ with dimension $(d+1)\times(d-1)$ has $2$ linearly independent left null vectors with high probability; let ${\mathbf{v}}_{0,1},{\mathbf{v}}_{0,2}\in{\mathbb{C}}^{1\times(d+1)}$ be the linearly independent left null vectors. Thus
${\mathbf{v}}_{0,i}{\mathbf{y}}^{\left[s\right]}$ $=\sum_{j=1}^{d-1}H^{\left[K+1,j\right]}{\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{\left[j\right]}+H^{\left[K+1,K+1\right]}v_{0,i}{\mathbf{x}}_{s}^{\left[K+1\right]}$
$+{\mathbf{v}}_{0,i}{\mathbf{z}}^{[{\mathsf{K}}+1]}.$
We then fix the values of ${\mathbf{x}}_{s}^{[{\mathsf{K}}+1]},{\mathbf{x}}_{s}^{[1]},\dots,{\mathbf{x}}_{s}^{[d-1]}$, the orthogonal sensing signals ${\mathbf{x}}_{i}^{[0]},i=\{1,2\}$ have been generated before. Let ${\mathbf{x}}_{i}^{[0]}=\begin{bmatrix}{\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]},{\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[1]},\ldots,{\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[d-1]}\end{bmatrix}^{T}$, we solve ${\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}$ with $d+1$ elements satisfying ${\mathbf{v}}_{0,i}{\mathbf{x}}_{s}^{[{\mathsf{K}}+1]}=x_{i,1}$, where $x_{i,1}$ is the first element of ${\mathbf{x}}_{i}^{[0]}$, and similarly for ${\mathbf{x}}_{s}^{[1]},\dots,{\mathbf{x}}_{s}^{[d-1]}$. Consequently, $2$ effective observations of the channel can be acquired, all transmission can achieve an sDoF equal to $\frac{2}{d+1}$.
Remark 3 (Connection of the proposed ISAC schemes based on TIM).
In this section, we propose bistatic ISAC schemes based on TIM, which leverage stable network connectivity patterns to eliminate interference without requiring exact CSI. We consider two canonical network topologies: the $({\mathsf{K}}+1,d)$-regular network and the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network, where the specific topology is determined by the underlying network structure.
  • These two networks give rise to two respective schemes. Although the connectivity patterns differ, both networks exhibit cyclic topological structures. Therefore, we adopt a cyclic coding strategy in both cases to exploit the network connectivity and align the interference signals at the receivers, thereby reducing the effective interference dimension.
  • However, the detailed construction of the linear coding schemes (based on subspace alignment) depends on the specific topology. In the $({\mathsf{K}}+1,d)$-regular network, transmitters are cooperative, and each receiver is connected to a sequence of $d$ consecutive transmitters in the topology. This allows each transmitter to send one symbol for its own target receiver and another for the receiver with the maximal topological distance, facilitating interference separation at the receivers.
    In contrast, in the $({\mathsf{K}}+1,U,D)$ neighboring antidotes network, each receiver is disconnected from a number of nearby interfering transmitters in the topology. This disconnection enables a cyclic coding design such that the signals from interfering transmitters are aligned into a lower-dimensional subspace at each receiver.